OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve (f + g)(3) and select the correct answer below: f(x) = 6x + 3 g(x) = x − 7 4 17 25 31

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

(f + g)(x) = 6x + 3 + x - 7

OpenStudy (welshfella):

simplify the above then replace the x by 3 and work it out

OpenStudy (anonymous):

I don't know how to simplify :P

OpenStudy (welshfella):

bring like terms together = 6x + x - 3 + 7 = 7x + 4 now replace the x by 3 and work it out this is your answer

OpenStudy (anonymous):

So it would be 17?

OpenStudy (welshfella):

7(3) + 4 = ? not 17

OpenStudy (anonymous):

25 :D?

OpenStudy (anonymous):

@welshfella